OpenStudy (anonymous):

(4/7)+3/(x-3)=(53/56) Solve Equation

OpenStudy (callisto):

\[\frac{4}{7}+\frac{3}{(x-3)}=\frac{53}{56}\]\[\frac{4(x-3)}{7(x-3)}+\frac{3(7)}{(x-3)(7)}=\frac{53}{56}\]\[\frac{4x-12+21}{7(x-3)}=\frac{53}{56}\]\[\frac{4x+9}{7(x-3)}=\frac{53}{56}\]\[(56)(4x+9)=53(7)(x-3)\]\[224x+504 = 371x - 1113\] Can you solve it now?